# Novel Application of Neutrinos to Evaluate U.S. Nuclear Weapons Performance

### v1

@@ Abstract
There is a growing realization that neutrinos can be used as a diagnostic tool to better understand the inner workings of a nuclear weapon. Robust estimates demonstrate that an Inverse Beta Decay (IBD) neutrino scintillation detector built at the Nevada Test Site of 1000-ton active target mass at a standoff distance of 500 m would detect thousands of neutrino events per kTe of nuclear yield 111we use the convention kTe for kilo-tons of explosive yield, while for neutrino detector mass we use ton.. This would provide less than 4% statistical error on measured neutrino rate and 5% error on neutrino energy. Extrapolating this to an error on the test device explosive yield requires knowledge from evaluated nuclear databases, non-equilibrium fission rates, and assumptions on internal neutron fluxes. Initial calculations demonstrate that prompt neutrino rates from a short pulse of 239Pu fission is about a factor of two less than that from a steady state assumption. As well, there are significant energy spectral differences as a function of time after the pulse that needs to be considered. These and other model dependencies will be discussed further in the paper. In the absence of nuclear weapons testing, many of the technical and theoretical challenges of a full nuclear test could be mitigated with a low cost smaller scale 20 ton fiducial mass IBD demonstration detector placed near a pulsed reactor. Potential reactors include the Texas A&M University (TAMU) TRIGA 1GW-10 millisecond pulsed facility or the Sandia Annular Core Research Reactor. The short duty cycle and repeatability of pulses would provide critical real environment testing and measurements which would be valuable for planning a possible real test shot in the future. Furthermore, the neutrino rate as a function of time data would provide unique constraints on fission databases and equilibrium assumptions. Finally, there are impactful science drivers such as sensitive searches for 1 eV2 sterile neutrinos and MeV scale axions.
@@ I Introduction
The application that we will be exploring in this paper is the use of a 1000-ton scale Inverse Beta Decay (IBD) neutrino detector to measure the nuclear weapon performance at a safe stand off distance. Situating the detector at a central accessible point allows it to be re-used for multiple shots. A 500 m standoff distance is sufficient to insulate the detector from shock waves but close enough to potentially detect thousands of neutrinos. The detector should be at least 10 m underground to shield from cosmic ray backgrounds. First, we perform robust rate calculations for neutrinos from a hypothetical yield to indicate if the concept is feasible. We have estimated the neutrino flux and neutrino event rates from a weapon test as a function of energy and time and have estimated fission fluxes by assuming a single prompt fission event which generates fragments with a known independent yield distribution. These products decay to a distribution of isotopes and the resulting neutrino spectrum is folded with known neutrino cross sections to produce rates for IBD. As well, a discussion of systematic errors associated with extrapolating the measured neutrino rates to nuclear detonation yield will be presented.
Second, the proven detection method for neutrinos from power reactors (U or Pu fission) is a large mass liquid scintillator detector (100-tons) that measures the IBD channel via [6]. The detection of the prompt positron followed by gamma rays from neutron capture provides a very powerful signal with excellent background rejection. The reconstructed positron interaction preserves the original neutrino energy and the event time to the nano-second level. Thus, the neutrino flux rate, energy, and time distribution can be accurately reconstructed. Initial straight forward rate estimates demonstrates that a large 1000-ton-scale detector 500 m from the nuclear test device will have significant event rates. We expect that neutrino rates can be measured with an uncertainty of four percent or better and energy at the 5% level. These robust detectors might be the key to detecting neutrinos with sufficient rate, energy, and time resolution to determine weapon yield and other parameters of interest to U.S. weapons design community.
Finally, we will discuss plans to test a small 25-ton IBD detector at a pulsed reactor facility. The aim is three fold; 1) test experimental issues operating a IBD detector at a pulsed fission source, 2) measure, for the first time, the neutrino time and energy distributions from a pulsed source, and 3) evaluate yield systematic errors from a pulsed non-equilibrium fission source using measured neutrino rates. This will be followed up by discussing possible fundamental and applied physics that can be accomplished by such a prototype detector at a pulsed reactor.
@@ II Estimated detected neutrino event rates from a potential nuclear weapons test
The fission processes produce neutrinos throughout the fission decay chain, which summed up total about 2.5 antineutrinos per U or Pu fission decay assuming non-equilibrium conditions of a fast fission pulse as described in Section III. For a steady state nuclear reactor this number is about 6 antineutrinos per fission [7]. This difference is significant and needs to be better understood and measured.
@@
Detailed flux calculations are readily available and show antineutrino spectra up to about 10 MeV [8]. Once produced, the neutrinos are emitted isotropically and most will pass through the Earth. However, with a sizable scintillation detector filled with hydrogen (such as mineral oil - CH2), some of the neutrinos can interact through the inverse beta decay mechanism, producing a time correlated positron and neutron in the final state that can be detected. Due to the threshold of 1.806 MeV for the reaction, the maximum energy the positron can deposit is 8 MeV in the form of scintillation and Cherenkov light in the detector medium. If the detector is doped with Gd, then the final state neutron thermalizes quickly and captures on the Gd within 10’s of micro-seconds producing up to 8 MeV of gamma-rays [9]. This time coincidence of the positron and neutron capture signals are unique and provide powerful rejection of random backgrounds.
@@
Decades of development of neutrino detectors to measure neutrino properties from nuclear reactors has produced a proven and robust design to detect neutrino IBD interactions [10, 11]. The basic detector design is a large volume (10 to 10,000 tons) of optically transparent scintillator material surrounded by large number of Photo-Multiplier Tubes (PMT’s) that detect the scintillation light from particle interactions. Fast digitizer electronics are used to read out the PMT’s which allows reconstruction of neutrino event rates, energy, position, and time of arrival. The next two sub-sections present neutrino IBD rate calculations for a mineral oil based detector instrumented with hundreds of PMT’s. We start with simple back of the envelope easy to understand calculations then move up to more complete calculations folding flux and cross sections with efficiency, ensuring consistency along the way.
II.0.1 Back of the Envelope
@@
The IBD neutrino event rate is given by . The flux average IBD cross section is roughly cm2 for a typical U or Pu fission decay neutrino flux [12]. The neutrino flux from a burst is estimated easily from energy considerations where 1 kTe of TNT J MeV of energy. With about 200 MeV/fission, this yields fissions. Including a factor of 2.5 for the average number of neutrinos per fission from a pulse of fast neutrons, as described in Section III [12], this gives about half a mole of neutrinos per 1 kTe. Assuming half a mole of neutrino fissions in one burst lasting 1000 seconds at a distance of 500 m, gives a neutrino flux average of . A distance of 500 m was considered a reasonable standoff distance for insuring detector integrity, assuming both the blast and detector are buried underground. The number of free in one metric ton of neutrino detector mineral oil is atoms. Putting these three numbers together gives an IBD neutrino rate of 0.41 events/kTe/ton at 500 m. This is assuming 100% efficiency. However, the reaction threshold for neutrino IBD reactions results in a 50% cross section efficiency above this energy. The detector reconstruction efficiency is estimated to be about 80% (see subsection IV.0.5), which results in a basic efficiency corrected neutrino rate of 0.16 events/kTe/ton at 500 m. This translates into thousands of detected neutrino events for a 1000-ton medium sized neutrino detector at a distance of 500 m from a plausible nuclear test. As well as rate, an IBD detector would be able to determine neutrino energy better than 5% resolution and event arrival time to nano-second level.
II.0.2 Spectrum Folding Calculations
A more detailed calculation of event rates is performed by folding the time dependent spectrum of neutrinos produced by fission products with the cross section for inverse decay. As an example, we take the energy spectrum of produced by fission products of 239Pu from Ref. [13] and Section III . We calculate the time dependent isotope yields after the prompt fission event, and show the time-integrated neutrino energy spectra in several windows of time after the fission pulse in the right panels of Fig. 2. The IBD energy threshold results in loss of sensitivity to the neutrino spectrum below the 1.8 MeV threshold, which is % of all generated neutrinos (hence, % cross section threshold efficiency). The total time-integrated neutrino spectrum, folded over the IBD cross section, is shown in the left panel of Fig. 2, while the right panel shows the spectrum for different time intervals. This shows that on average higher energy neutrinos come in the earliest time bins after the prompt fission pulse.
A detector with 1000-tons of active mass would have a radius of 6 m. Such an instrument placed 500 m from a test device subtends a fraction of of the solid angle of the neutrinos emitted by the device. A typical Gd-doped mineral oil produced by Eljen, EJ-335, [14], has a hydrogen density of H atoms per cubic centimeter. This gives an event rate of 0.16 events/kTe/ton, similar to the back-of-the-envelope calculation quoted above. Uncertainties associated with the IBD cross section can be evaluated in future work, but are expected to be less than . Due to the threshold and rising neutrino cross section, the detectable neutrino energy spectrum shown in Figure 2 peaks around 4 MeV, with an endpoint energy near 8 MeV. Figure 3 shows the neutrino production rate from a short fission burst, with approximately 30% of the total neutrinos arriving before 10 seconds, though there is a long tail up to 100’s of seconds.
@@ III.0.1 Extracting fission yields from the emitted neutrinos
Reactor thermal fission neutrino spectra are known at the peak of the spectrum when folded with the IBD cross section to an accuracy of approximately 3%, however the uncertainties are energy dependent and at the largest energies (above approximately 8 MeV) the uncertainty is approximately 20% [15]. In a nuclear explosion, the neutron energies are considerably higher than in a reactor where the neutrons inducing fission in a nuclear explosion are fast neutrons (with energies of 0.4-2.0 MeV) and 14 MeV neutrons. Extrapolating from measured reactor thermal spectra to those expected for weapons-relevant neutron energies, with reasonable accuracy, is made possible by the very detailed systematic studies of the energy-dependence of both the independent and cumulative fission product yields (FPYs) that have been carried out at Los Alamos and elsewhere, in a joint long-term experimental and theoretical program [16].
@@
| 0.0-0.1 s | 0.1-1.0 s | 1.0-10.0 s | 10.0-100.0 s | 100.0-1000.0 s | Total | |
| (Fast) | 0.020 | 0.145 | 0.572 | 0.890 | 0.864 | 2.491 |
| (14 MeV) | 0.018 | 0.129 | 0.479 | 0.697 | 0.707 | 2.03 |
@@
An IBD detector is ultimately sensitive only to neutrinos with energies above the IBD threshold of approximately 1.8 MeV. The total number of events seen by the detector between times and is proportional to the spectrally averaged IBD cross section .
@@
Assuming no variation in the total neutrino flux across the volume of the detector which is situated a distance from the prompt fission event (e.g. a nuclear explosion), the total number of detected IBD events in a window of time () is then
@@
Here, is the total number of protons in the detector and is the total number of fissions which happened in the pulse. Thus, is the effective cross sectional area that the IBD detector presents to the antineutrinos from the explosion. Because the antineutrino spectrum evolves in time, will also vary through time. We have tabulated it for the same windows in time as Table 1 and present it in Table 2.
@@
There are two opposing effects at work in the evolution of . Because -decay is a weak decay process, the beta decay of fission fragments occurs on a much slower timescale than the fission event itself and thus there are relatively few antineutrinos emitted at early times. However, the early-time antineutrinos carry a higher energy on average than those at later times. Because the IBD cross section above the reaction energy threshold is an increasing function of energy, rises to a peak value after approximately seconds and then falls as both the antineutrino emission rate and average antineutrino energy fall. We have tabulated the falling average antineutrino energy in Table 3 for the same time windows. We have not accounted for the effect of antineutrino flavor oscillations in this analysis, as the oscillation probability depends on the distance which we take as a free parameter in this discussion. The values presented here are therefore an upper bound on the detector response, and a more careful treatment applicable to a specific detector would need to take the disappearance of through flavor oscillations into account by weighting the integrand of eq. 1 by the energy and distance dependent survival probability.
| 0.0-0.1 s | 0.1-1.0 s | 1.0-10.0 s | 10.0-100.0 s | 100.0-1000.0 s | Total | |
| (Fast) | 0.077 | 0.515 | 1.603 | 1.375 | 0.666 | 4.236 |
| (14 MeV) | 0.072 | 0.460 | 1.288 | 0.991 | 0.540 | 3.351 |
| 0.0-0.1 s | 0.1-1.0 s | 1.0-10.0 s | 10.0-100.0 s | 100.0-1000.0 s | |
| (Fast) | 2.842 | 2.765 | 2.522 | 2.042 | 1.587 |
| (14 MeV) | 2.871 | 2.773 | 2.495 | 1.988 | 1.581 |
@@
In extracting the number of fissions from a neutrino measurement, we first consider the scenario in which there is no independent measurement/knowledge of the 14 MeV neutron yield, except from a simulation prediction of unknown accuracy. We also assume that the sole actinide fissioning is 239Pu. The predicted measured neutrino spectrum in Fig. 2 is the sum arising from fast and 14 MeV neutron-induced fission. Clearly any uncertainty in neutron fluxes will result in nuclear explosion yield errors when extrapolating from measured neutrino fluxes.

### v2

@@ Abstract
There is a growing realization that neutrinos can be used as a diagnostic tool to better understand the inner workings of a nuclear weapon. Robust estimates demonstrate that an Inverse Beta Decay (IBD) neutrino scintillation detector built at the Nevada Test Site of 1000-ton active target mass at a standoff distance of 500 m would detect thousands of antineutrino events per nuclear test. This would provide less than 4% statistical error on measured antineutrino rate and 5% error on antineutrino energy. Extrapolating this to an error on the test device explosive yield requires knowledge from evaluated nuclear databases, non-equilibrium fission rates, and assumptions on internal neutron fluxes. Initial calculations demonstrate that the total number of neutrinos emitted per fission in the first s after a short pulse of 239Pu fission is about a factor of two less than that from Pu fissioning under steady state conditions. As well, there are significant energy spectral differences as a function of time after the pulse that must be considered. These and other model dependencies will be discussed in the paper. In the absence of nuclear weapons testing, many of the technical and theoretical challenges of a full nuclear test could be mitigated with a low cost smaller scale 20 ton fiducial mass IBD demonstration detector placed near a pulsed reactor. Potential reactors include the Texas A&M University (TAMU) TRIGA 1GW-10 millisecond pulsed facility or the Sandia Annular Core Research Reactor. The short duty cycle and repeatability of pulses would provide critical real environment testing and measurements which would be valuable for planning a possible real test shot in the future. Furthermore, the antineutrino rate as a function of time data would provide unique constraints on fission databases and model assumptions. Finally, there are impactful science drivers such as sensitive searches for 1 eV2 sterile neutrinos and MeV scale axions.
@@ I Introduction
The application that we will be exploring in this paper is the use of a 1000-ton scale Inverse Beta Decay (IBD) neutrino detector to measure the nuclear weapon performance at a safe stand off distance. Situating the detector at a central accessible point allows it to be re-used for multiple shots. A 500 m standoff distance is sufficient to insulate the detector from shock waves but close enough to potentially detect thousands of antineutrinos from a test. The detector should be at least 10 m underground to shield from cosmic ray backgrounds. First, we perform robust rate calculations for antineutrinos from a hypothetical nuclear yield to indicate if the concept is feasible. We estimate the antineutrino flux and IBD event rates from a weapon test as a function of energy and time by assuming a single prompt fission event which generates fragments with a known independent yield distribution. These products decay to a distribution of stable (or sufficiently long lived) isotopes and the resulting antineutrino spectrum is folded with known antineutrino cross sections to produce rates for IBD interactions. A discussion of systematic errors associated with extrapolating the measured neutrino rates to nuclear detonation yield will be presented.
The proven detection method for neutrinos from power reactors (U or Pu fission) is a large mass liquid scintillator detector (100-tons) that measures the IBD channel via [6]. The detection of the prompt positron followed by gamma rays from neutron capture provides a very powerful signal with excellent background rejection. The reconstructed positron interaction preserves the original antineutrino energy and the event time to the nanosecond level. Thus, the antineutrino flux rate, energy, and time distribution can be accurately reconstructed. Initial straight forward rate estimates demonstrate that a large 1000-ton-scale detector 500 m from the nuclear test device will have significant event rates per kTe of nuclear yield 111Throughout this work we use the convention kTe for kilo-tons of explosive yield, while for neutrino detector mass we use ton.. We expect that antineutrino rates can be measured with an uncertainty of four percent or better and energy at the 5% level. These robust detectors might be the key to detecting neutrinos with sufficient rate, energy, and time resolution to determine weapon yield and other parameters of interest to the U.S. weapons design community.
Finally, we discuss plans to test a small 25-ton IBD detector at a pulsed reactor facility. The aims are three fold: 1) to test experimental issues operating an IBD detector at a pulsed fission source, 2) to measure, for the first time, the antineutrino time and energy distributions from a pulsed source, and 3) to evaluate yield systematic errors from a pulsed non-equilibrium fission source using measured neutrino rates. This will be followed by discussing possible fundamental and applied physics that can be accomplished by such a prototype detector at a pulsed reactor.
@@ II Estimated detected neutrino event rates from a potential nuclear weapons test
The fission processes produce neutrinos throughout the fission decay chain, which summed up total about 2.5 antineutrinos per U or Pu fission decay in the first seconds after a fast fission pulse as described in Section III. For a steady state nuclear reactor this number is about 5 antineutrinos per fission (see table 1).
@@
Detailed flux calculations are readily available and show antineutrino spectra up to about 10 MeV [8]. Once produced, the antineutrinos are emitted isotropically and will pass through the ground. However, with a sizable scintillation detector filled with hydrogen (such as mineral oil - CH2), some of the antineutrinos can interact through the inverse beta decay mechanism, producing a time correlated positron and neutron in the final state that can be detected. Due to the threshold of 1.806 MeV for the reaction, the maximum energy the positron can deposit is 8 MeV in the form of scintillation light in the detector medium. If the detector is doped with Gd, then the final state neutron thermalizes quickly and captures on the Gd within 10’s of micro-seconds producing up to 8 MeV of gamma-rays [9]. The time coincidence of the positron and neutron capture signals are unique and provide powerful rejection of random backgrounds.
@@
Decades of development of neutrino detectors to measure neutrino properties from nuclear reactors has produced a proven and robust design to detect antineutrino IBD interactions [10, 11]. The basic detector design is a large volume (10 to 10,000 tons) of optically transparent scintillator material surrounded by a large number of Photo-Multiplier Tubes (PMT’s) that detect scintillation light from particle interactions. Fast digitizer electronics are used to read out the PMT’s which allows reconstruction of antineutrino event rates, energy, position, and time of arrival. The next subsection presents a simple antineutrino IBD rate calculation for a mineral oil based detector instrumented with hundreds of PMT’s. We start with back of the envelope calculations then move up to more complete calculations folding flux and cross sections with efficiency, ensuring consistency along the way.
II.0.1 Estimating the measurable antineutrino signal
@@
We can estimate the number of antineutrino events that a typical nuclear fission explosion would generate per kiloton of TNT explosive equivalent nuclear energy (kTe) released in an IBD detector of 1000-ton typical fiducial mass at a 500m typical standoff distance. First, the IBD event rate is given by where is the number of free hydrogen targets in the detector. The typical flux averaged cross section is roughly cm2 for a typical U or Pu fission decay neutrino flux [12]. The antineutrino flux from a burst is estimated easily from energy considerations where 1 kTe of TNT J MeV of energy. With about 200 MeV/fission [13], this corresponds to fissions. Including a factor of 2.5 for the average number of antineutrinos per fission emitted in the first s after a pulse of fissions induced by “fast” neutrons (e.g. neutrons emitted during fission carrying energies between 0.4-2 MeV), as described in Section III, this gives about half a mole of antineutrinos per 1 kTe. Using this, the flux at m from the explosion is estimated to be cm-2. A distance of 500 m was considered a reasonable standoff distance for ensuring detector integrity, assuming that both the blast and detector are buried underground. The number of free in one metric ton of neutrino detector mineral oil is atoms. Putting these three numbers together gives an IBD neutrino rate of 0.41 events/kTe/ton at 500 m. This is assuming 100% efficiency. However, the reaction threshold for neutrino IBD reactions results in a 50% cross section efficiency above this energy. The detector reconstruction efficiency is estimated to be about 80% (see subsection IV.0.5), which results in a basic efficiency corrected neutrino rate of 0.16 events/kTe/ton at 500 m. This translates into thousands of detected neutrino events for a 1000-ton neutrino detector at a distance of 500 m from a plausible nuclear test [14]. As well as rate, an IBD detector would be able to determine neutrino energy better than 5% resolution and event arrival time to nanosecond level.
@@ III Measurements, Extraction of Nuclear Test Yield, Systematic Errors
In a reactor, thermal neutrons induce fissions of the actinide fuel resulting in a nearly steady state (herein abreviated as “S.S.”) emission of antineutrinos. The weighted product of this antineutrino spectrum with the IBD cross section is known to an accuracy of approximately 3% at its peak value, but the uncertainties are energy dependent. At the highest energies (above 8 MeV) the uncertainty is approximately 20% [15]. The typical energies of neutrons which induce fission in a nuclear explosion are considerably higher than in a reactor environment. In a weapon environment, the fissioning neutrons are dominantly a combination of “fast” (e.g. those emitted from a fissioning nucleus and carrying an energy of 0.4-2.0 MeV) and 14 MeV in energy.
Extrapolating from measured reactor thermal spectra to those expected for weapons-relevant neutron energies, with reasonable accuracy, is made possible by the very detailed systematic studies of the energy-dependence of both the independent and cumulative fission product yields (FPY’s) that have been carried out at Los Alamos and elsewhere, in a joint long-term experimental and theoretical program [16].
@@ III.0.1 Extracting fission yields from the emitted neutrinos
| 239Pu | 0.0-0.1 s | 0.1-1.0 s | 1.0-10.0 s | 10.0-100.0 s | 100.0-1000.0 s | Total | S.S. |
| (Fast) | 0.020 | 0.145 | 0.572 | 0.890 | 0.864 | 2.49 | 4.34 |
| (14 MeV) | 0.018 | 0.129 | 0.479 | 0.697 | 0.707 | 2.03 | 3.63 |
| Fast / 14 MeV | 1.11 | 1.12 | 1.19 | 1.27 | 1.22 | 1.23 | 1.19 |
| 235U | |||||||
| (Fast) | 0.032 | 0.219 | 0.774 | 1.013 | 0.856 | 2.89 | 4.80 |
@@
An IBD detector is ultimately sensitive only to neutrinos with energies above the IBD threshold of approximately 1.8 MeV. The total number of events seen by the detector between times and is proportional to the spectrally averaged IBD cross section . Explicitly,
@@
Assuming no variation in the total neutrino flux across the volume of the detector which is situated a distance from the prompt fission event (e.g. a nuclear explosion), the total number of detectable IBD events in a window of time () is then
@@
Here, is the total number of protons in the detector and is the total number of fissions which happened in the pulse. Thus, is the (time dependent) cross sectional area that the IBD detector presents to the evolving spectrum of antineutrinos in that window of time after the explosion. We have tabulated it for the same windows in time as table 1 and present it in table 2.
There are two opposing effects at work in the evolution of . Because -decay is a weak decay process, the -decay of fission fragments occurs on a much slower timescale than the fission event itself and thus there are relatively fewer antineutrinos emitted at early times. However, the early-time antineutrinos carry a higher energy on average than those at later times. In the time between and , the average antineutrino energy is given by
@@
Because the IBD cross section above the reaction energy threshold is an increasing function of energy, rises to a peak value after approximately seconds and then falls as both the antineutrino emission rate and average antineutrino energy fall. We have tabulated the falling time-averaged antineutrino average energy (eq. 3) in Table 3 for the same time windows. We have not accounted for the effect of antineutrino flavor oscillations in this analysis, as the oscillation probability depends on the distance which we take as a free parameter in this discussion. The values presented here are therefore an upper bound on the detector response, and a more careful treatment applicable to a specific detector would need to take the disappearance of through flavor oscillations into account by weighting the integrand of eq. 1 by the energy and distance dependent survival probability.
| 239Pu | 0.0-0.1 s | 0.1-1.0 s | 1.0-10.0 s | 10.0-100.0 s | 100.0-1000.0 s | Total | S.S. |
| (Fast) | 0.077 | 0.515 | 1.603 | 1.375 | 0.666 | 4.236 | 4.65 |
| (14 MeV) | 0.072 | 0.460 | 1.288 | 0.991 | 0.540 | 3.351 | 3.79 |
| 235U | |||||||
| (Fast) | 0.146 | 0.911 | 2.574 | 1.838 | 0.782 | 6.25 | 6.77 |
| 239Pu | 0.0-0.1 s | 0.1-1.0 s | 1.0-10.0 s | 10.0-100.0 s | 100.0-1000.0 s | S.S. |
| (Fast) | 2.842 | 2.765 | 2.522 | 2.042 | 1.587 | 1.61 |
| (14 MeV) | 2.871 | 2.773 | 2.495 | 1.988 | 1.581 | 1.59 |
| 235U | ||||||
| (Fast) | 3.020 | 2.914 | 2.651 | 2.118 | 1.633 | 1.748 |
@@
In extracting the number of fissions from a neutrino measurement, we first consider the scenario in which there is no independent measurement/knowledge of the 14 MeV neutron yield, except from a simulation prediction of unknown accuracy. We also assume that the sole actinide fissioning is 239Pu. Clearly any uncertainty in neutron fluxes will result in nuclear explosion yield errors when extrapolating from measured neutrino fluxes.